# Binarization-Aware Adjuster: Bridging Continuous Optimization and Binary Inference in Edge Detection

### v1

@@ V Conclusion
This paper tackles a fundamental yet underexplored issue in image edge detection (ED): the misalignment between the continuous-valued outputs optimized during training and the binary predictions required at inference and evaluation. This discrepancy, often overlooked due to the discontinuity of binarization, undermines the alignment between training objectives and final task performance.
@@
To address this, we propose a principled binarization adjustment mechanism that incorporates binarization awareness into the training process. Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each pixel’s contribution to the loss based on its prediction correctness and confidence. It downweights confidently correct predictions since further optimizing them has little effect on the binary output, while emphasizing ambiguous or misclassified regions that are more critical to the final binarization decision. The resulting adjusted loss function from the BAA generalizes traditional loss formulations and embeds binarization behavior directly into the optimization process. In addition, we introduce a self-adaptive training procedure to dynamically estimate an appropriate binarization threshold for the BAA, further aligning training dynamics with inference behavior.
Extensive experiments across diverse models and datasets validate the effectiveness and robustness of the proposed method. Beyond ED, our BAA approach can also provide a theoretical method for bridging the gap between continuous optimization and discrete evaluation, offering a promising direction for broader applications in structured prediction tasks such as segmentation and classification.

### v2

@@ 6 Conclusion
This paper tackles a fundamental yet underexplored issue occurring in discrete decision machine learning tasks: the misalignment between the continuous-valued outputs optimized during training and the discrete predictions required at inference and evaluation. This discrepancy, often overlooked due to the discontinuity of discretization, undermines the alignment between training objectives and final task performance.
@@
To address this, we propose a principled binarization adjustment mechanism that incorporates binarization awareness into the training process. Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each element’s contribution to the loss based on its prediction correctness and confidence. It downweights confidently correct predictions since further optimizing them has little effect on the binary output, while emphasizing ambiguous or misclassified regions that are more critical to the final binarization decision. The resulting adjusted loss function from the BAA generalizes traditional loss formulations and embeds binarization behavior directly into the optimization process. In addition, we introduce a self-adaptive training procedure to dynamically estimate an appropriate binarization threshold for the BAA, further aligning training dynamics with inference behavior. Our method is applied to the edge detection (ED) task, and extensive experiments across representative models and datasets validate its effectiveness and robustness.
In summary, the BAA introduces a theoretically grounded mechanism for bridging continuous optimization and discrete evaluation, with application to the ED task, demonstrating the effectiveness experimentally. And beyond ED task and binarization decision tasks, the principle can inform related discrete structured-prediction tasks, offering a promising direction for broader applications.